OpenStudy (sonyalee77):

PLEASE HELP WILL GIVE MEDAL!!!!! Graph the inequality on a number line:-5[y+4]>=3[y-4]

OpenStudy (anonymous):

-5(y+4)>=3(y-4) -5y-20+12 >=3y-12+12 -5y-8+5y >=3y+5y -8>=8y y <=-1

OpenStudy (anonymous):

can u graph y <=-1?